Information: Suppose the firm has 500 employees each working a standard workweek of 40 hours. The...

  1. Information: Suppose the firm has 500 employees each working a standard workweek of 40 hours. The typical worker earns a real wage of 10 per hour and receives another 200 in weekly benefits and training, 150 of which are quasi-fixed costs to the firm.

    The marginal expense an employer faces when increasing its average workweek by an hour (MEH) is approximately the sum of wage and variable benefit costs (on a per-hour basis) multiplied by he number of employees. Assuming no overtime wage premium, an approximate value for MEH would be

    11.25

    60

    5,625

    30,000

10 points   

QUESTION 2

  1. Suppose the firm has 500 employees each working a standard workweek of 40 hours. The typical worker earns a real wage of 10 per hour and receives another 200 in weekly benefits and training, 150 of which are quasi-fixed costs to the firm.Suppose the marginal product of an additional worker is 50 and the marginal product associatedwith expanding the workweek by 1 hour is 625. Using the MEH and MEM values computed in Questions 6 and 7, employ the rule for cost minimization to find the firm’s optimal adjustment of its employment/hours mix.

    Increase M and decrease H.

    Increase H and decrease M

    Increase both H and M.

    M and H should not change.

10 points   

QUESTION 3

  1. Consider again the law that would require employers to provide the same package of nonwage benefits offered to full-time employees to part-time employees on a prorated basis. (For employees working 25 percent of the normal workweek, employers would pay for 25 percent of the benefit package offered to full-time employees, and so on.) Also, assume most part-time workers do not currently receive benefits. Holding output and capital constant, what effect would there be on the firm’s mix of part-time and full-time workers?

    Full-time workers would increase and part-time workers would decrease since the marginal cost of part-time workers is now higher.

    Full-time workers would decrease and part-time workers would decrease since the marginal cost of part-time workers is now higher.

    There would be no change in the mix of part-time and full-time workers since part-time workers are still cheaper than full-time workers.

    Part-time employment would actually increase since more people would now want to work part-time.

10 points   

QUESTION 4

  1. Which of the following is not a quasi-fixed cost?

    The hourly wage the firm pays.

    Costs associated with a defined benefit pension .

    Costs associated with a defined contribution pension.

    Employer contributions to Social Security.
Answer #1

Question 1

The answer is 5625.

MEM - the marginal expense of hiring an added worker

MEH - the marginal expense of hiring current workers for an extra hour

Variable benefits per hour = (200-150)/40 = 50/40

MEH = (wage + variable benefits per hour) × number of workers = [10 + (50/40)] 500 = 5,625.